2 cars are travelling on saparate lanes on a 2 lane freeway.how long does it take car 2 travelling at 100km heitz per second to overtake car 1 travelling at 80km heitz per second if the distance between them is 150m
Answer by Alan3354(69443)

2 cars are travelling on saparate lanes on a 2 lane freeway.how long does it take car 2 travelling at 100km heitz per second to overtake car 1 travelling at 80km heitz per second if the distance between them is 150m
--------------
What is "heitz?"
The difference in speed is 20 km/sec
d = r*t
t = d/r = 0.15 km/20 km/sec
t = 0.0075 seconds or 7.5 milliseconds
----------
Those are fast cars.